The sum of anne's amd fred's ages are 28. in 7 years anne will be twice as old as fred. How old are the now?
Answer by TutorDelphia(193)

a+f=28
In 7 years a will be a+7 and f will be f+7
a+7=2(f+7)
a+7=2f+14
a=2f+7
sub in 2f+7 for a into my first equation and I get
2f+7+f=28
3f=21
f=7
Plug in 7 for f
a+7=28
a=21